OpenStudy (anonymous):

sqrt(8 − 2x) = x − 4

hartnn (hartnn):

square both sides, what u get ?

OpenStudy (anonymous):

8-2x=x-4

hartnn (hartnn):

the right side also gets squared 8-2x = (x-4)^2 now expand right side.......

OpenStudy (anonymous):

8-2x=x^2-16?

hartnn (hartnn):

(x-a)^2 = x^2-2ax+a^2

hartnn (hartnn):

so, 8-2x = x^2-8x+16 now bring everything on one side

OpenStudy (anonymous):

24-10x=x^2?

hartnn (hartnn):

umm, nopes \(8-2x = x^2-8x+16 \\ x^2-8x+2x+16-8=0 \\ x^2-6x+8=0\) can u solve this quadratic ?

OpenStudy (anonymous):

Would it be lame if i said i didn't know how xD

hartnn (hartnn):

no, its okay. split the middle term \(x^2-4x-2x+8=0\)

hartnn (hartnn):

\(x(x-4)-2(x-4)=0\) can u solve ahead ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

hartnn (hartnn):

what u get?

OpenStudy (anonymous):

x^2-4x-2x-8=0

OpenStudy (anonymous):

then combine like terms?

hartnn (hartnn):

why u going reverse way? u will get your question back :P \(x(x-4)-2(x-4)=0 \\(x-4)(x-2)=0 \\ x=4 \:or\:x=2\) got this ?

OpenStudy (anonymous):

Idk im dumb xD lol thank you!

hartnn (hartnn):

welcome ^_^